# MMSummary: Multimodal Summary Generation for Fetal Ultrasound Video

### v1

@@ 2 Method
This stage detects a concise subset of representative and diverse keyframes that are distinct from each other, summarizing the useful anatomical content in the video. As in Fig. 2 (left side), frames of the video are passed through a pre-trained image feature extractor ($f_{I}(\cdot)$, BiomedCLIP ViT-B [33]) followed by a learnable linear layer, yielding a sequence of feature embeddings. These embeddings are added to vanilla sine-cosine temporal position embeddings, and fed into a 4-layer video transformer $f_{V}(\cdot)$ (each containing a multi-head self-attention and MLP), outputting the probability of each frame being a keyframe $p^{\prime}\in[0,1]^{T}$. To optimize the learnable linear layer and video transformer, binary cross-entropy loss $\mathcal{L}_{BCE}(p^{\prime},y)=-y\cdot log(p^{\prime})-(1-y)\cdot log(1-p^{\prime})$ is computed w.r.t. the keyframe label $y$.
@@
To mitigate redundancy in detected frames, we further propose an diverse keyframe detection algorithm (Fig. 4 left), which iteratively selects keyframes to ensure exclusiveness in the resulting subset. In each iteration, the frame with the highest probability in the video is selected as the keyframe, and then we mask out other frames that share similar features with the selected keyframe (i.e., have a similarity above $\tau$). This process continues iteratively until the highest probability falls below $\tau^{\prime}$. Finally, we obtain a concise set of keyframes.
It is worth noting that choosing good values for $\tau$ and $\tau^{\prime}$ is crucial, as shown in Fig. 5. A too large $\tau$ cannot thoroughly filter out similar frames, leading to frame redundancy. Conversely, a too small $\tau$ may result in a risk of information loss as some representative frames are discarded. As for $\tau^{\prime}$, a too small value may lead to the inclusion of non-informative frames, while choosing a too large value may result in information loss due to discarding potentially useful frames. Hence, well chosen $\tau$ and $\tau^{\prime}$ values can ensure the selected keyframes effectively capture the essential information while minimizing redundancy and information loss.
@@ 3 Experiments
Dataset. The dataset utilized in this study is from [anonymized project], with approval from [anonymized Ethics Committee]. Clinical fetal US scans were conducted using a GE Voluson E8 scanner. The videos were captured at 30 Hz and downsampled to 5 Hz. Our study included 178 second-trimester routine examination videos conducted by 3 newly qualified (less than two years of experience) and 1 expert (14 years of experience) sonographers. These videos were randomly split into training (145), validation (11), and test (22) sets. The ground-truth keyframes containing meaningful anatomical structures were obtained from screenshots captured by sonographers during clinical scanning. This resulted in 4596, 329, 679 keyframes in the three sets, of which 3481, 254, 531 frames had caption labels describing anatomies. For biometry estimation, we annotated bounding boxes to identify regions of interest and used MedSAM [18] for mask label inference. Due to the poor quality of generated mask labels for the cerebellum and femur classes, manual reannotation was performed, resulting in 574, 43, 75 mask labels. Ground-truth biometric parameters were manually annotated by sonographers during the scanning process.
Evaluation metrics. In keyframe detection (stage 1), we used four metrics to evaluate the performance of predicted keyframes $p$ compared with the ground truth $y$, as detailed in the right part of Fig. 4. 1) ‘Cosine simi.’ and 2) ‘absolute time err.’ measure keyframe feature similarity and absolute timestamp error in a set-to-set manner. As some predicted keyframes exhibiting large time errors share the same appearance as the ground truth, 3) ‘correct time err.’ measures the absolute keyframe timestamp error only when the maximum feature similarity is below a certain threshold (e.g., 0.96). 4) ‘Keyframe num. err.’ indicates the absolute error in the number of predicted keyframes. In the keyframe captioning (stage 2), two conventional metrics ‘BLEU’ and ‘ROUGE-L’ were utilized, following [1]. For the final stage 3, absolute error and relative error of AC, HC, BPD, Cereb, FL biometry w.r.t. the clinical measurement were computed.

### v2

@@ 2 Method
This stage detects a concise subset of representative and diverse keyframes that are distinct from each other, summarizing the useful anatomical content in the video. As in Figure 2 (left side), frames of the video are passed through a pre-trained image feature extractor ($f_{I}(\cdot)$, BiomedCLIP ViT-B [34]) followed by a learnable linear layer, yielding a sequence of feature embeddings. These embeddings are added to vanilla sine-cosine temporal position embeddings, and fed into a 4-layer video transformer $f_{V}(\cdot)$ (each containing a multi-head self-attention and MLP), outputting the probability of each frame being a keyframe $p^{\prime}\in[0,1]^{T}$. To optimize the learnable linear layer and video transformer, binary cross-entropy loss $\mathcal{L}_{BCE}(p^{\prime},y)=-y\cdot log(p^{\prime})-(1-y)\cdot log(1-p^{\prime})$ is computed w.r.t. the keyframe label $y$.
@@
To mitigate redundancy in detected frames, we further propose an diverse keyframe detection algorithm (supp. Fig. 4), which iteratively selects keyframes to ensure exclusiveness in the resulting subset. In each iteration, the frame with the highest probability in the video is selected as the keyframe, and then we mask out other frames that share similar features with the selected keyframe (i.e., have a similarity above $\tau$). This process continues iteratively until the highest probability falls below $\tau^{\prime}$. Finally, we obtain a concise set of keyframes.
It is worth noting that choosing good values for $\tau$ and $\tau^{\prime}$ is crucial (supp. Fig. 5). A too large $\tau$ cannot thoroughly filter out similar frames, leading to frame redundancy. Conversely, a too small $\tau$ may result in a risk of information loss as some representative frames are discarded. As for $\tau^{\prime}$, a too small value may lead to the inclusion of non-informative frames, while choosing a too large value may result in information loss due to discarding potentially useful frames. Hence, well chosen $\tau$ and $\tau^{\prime}$ values can ensure the selected keyframes effectively capture the essential information while minimizing redundancy and information loss.
@@ 3 Experiments
Dataset. The dataset utilized in this study is from PULSE (Perception Ultrasound by Learning Sonographic Experience) [7], with approval fromUK Research Ethics Committee. Clinical fetal US scans were conducted using a GE Voluson E8 scanner. The videos were captured at 30 Hz and downsampled to 5 Hz. Our study included 178 second-trimester routine examination videos conducted by 3 newly qualified (less than two years of experience) and 1 expert (14 years of experience) sonographers. These videos were randomly split into training (145), validation (11), and test (22) sets. The ground-truth keyframes containing meaningful anatomical structures were obtained from screenshots captured by sonographers during clinical scanning. This resulted in 4596, 329, 679 keyframes in the three sets, of which 3481, 254, 531 frames had caption labels describing anatomies. For biometry estimation, we annotated bounding boxes to identify regions of interest and used MedSAM [19] for mask label inference. Due to the poor quality of generated mask labels for the cerebellum and femur classes, manual reannotation was performed, resulting in 574, 43, 75 mask labels. Ground-truth biometric parameters were manually annotated by sonographers during the scanning process.
Evaluation metrics. In keyframe detection (stage 1), we used four metrics to evaluate the performance of predicted keyframes $p$ compared with the ground truth $y$, as detailed in Fig. 4 of supplementary. 1) ‘Cosine simi.’ and 2) ‘absolute time err.’ measure keyframe feature similarity and absolute timestamp error in a set-to-set manner. As some predicted keyframes exhibiting large time errors share the same appearance as the ground truth, 3) ‘correct time err.’ measures the absolute keyframe timestamp error only when the maximum feature similarity is below a certain threshold (e.g., 0.96). 4) ‘Keyframe num. err.’ indicates the absolute error in the number of predicted keyframes. In the keyframe captioning (stage 2), two conventional metrics ‘BLEU’ and ‘ROUGE-L’ were utilized, following [1]. For the final stage 3, absolute error and relative error of AC, HC, BPD, Cereb, FL biometry w.r.t. the clinical measurement were computed.